3. A customer support center for Bell Computer Company (BMC) receives an average of 2.5 phone calls every 5 minutes. Assume that the number of calls received follows Poisson distribution with λ = 2.5 in answering the following questions

  1. What is the probability that no calls will arrive during the next five minutes?
  2. What is the probability that 5 calls will arrive during the next 5 minutes?
  3. What is the probability that at least 3 calls will arrive during the next five minutes?
  4. What is the probability that at least 3 calls will arrive during the next 10 minutes?
  5. What is the probability that no more than 3 calls will arrive during the next 10 minutes?
Answer #1


Part a)

Part b)

Part c)



P ( X >= 3 ) = 1 - P ( X <= 2 )

P ( X >= 3 ) = 0.4562

Part d)



P ( X >= 3 ) = 1 - P ( X <= 2 )

P ( X >= 3 ) = 0.8754

Part e)



P ( X <= 3 ) = 0.2650